# silent-chen.github.io/DGE

### v1

@@ 4.2 Consistent Multi-view Editing with Epipolar Constraint
In order to achieve multi-view consistent edits, we consider the views $\mathcal{I}$ rendered from $\mathcal{G}$ as if they were arising from a relatively smooth camera trajectory and thus form a video. In this way, we reduce consistent multi-view editing to video editing.
While there is no off-the-shelf fast and video editor, we can take advantage of recent methods that extend image generators to video generators without additional training [24, 106, 44, 8, 80]. By applying them to image editors instead, we can obtain corresponding video editors.
@@
Let $\mathcal{E}$ be a diffusion-based image editor that utilizes self-attention (e.g., SD-based). Based on the above insight, we propose to extend $\mathcal{E}$ with a modified, spatio-temporal version of self-attention [99, 24], which jointly considers all frames in $\mathcal{I}$, effectively turning the single-image editor into a multi-image editor $\mathcal{E_{\operatorname{mv}}}$. Specifically, given queries $\{Q_{t}\}_{t=1}^{T}$, keys $\{K_{t}\}_{t=1}^{T}$, and values $\{V_{t}\}_{t=1}^{T}$ for each viewpoint $t$, the spatio-temporal attention block ensures that each frame attends to every other frame and is given by:
@@
$\operatorname{STAttn}(Q,K,t)=\operatorname{Softmax}\left(\frac{Q_{t}[\,K_{1},\dots,K_{T}]\ }{\sqrt{d}}\right),$ (5)
where $d$ is the embedding dimension of keys and queries. The corresponding output features for frame $t$ are then computed as $\Phi_{t}=\operatorname{STAttn}(Q,K,t)\cdot[\,V_{1},\dots,V_{T}]$ (at each layer).
As the underlying image editor $\mathcal{E}$, we utilize IP2P [4], which is a common choice in the literature. Following the framework of [23], each denoising iteration of the editing process consists of two steps.
The first step is keyframe editing. Because attention can be computationally expensive for longer sequences, $\mathcal{E_{\operatorname{mv}}}$ is only applied to a number of keyframes. In practice, we first sort the cameras given their pose matrix to have smooth views. Then, we randomly subsample the video (viewpoint) sequence at each denoising iteration. We detail the sorting process in the supplementary. This step results in roughly consistent edits across all keyframes.
@@
Figure 1: Overview. Our method is divided into two main parts: multi-view consistent editing and direct 3D fitting. In the multi-view editing stage, keyframes are randomly selected and fed to the editing diffusion network to extract features. In the editing process of other frames, the features of keyframes are injected into the diffusion network with the proposed epipolar constraint; only features in the red boxes are considered for correspondence matching to have a consistent appearance.
The second step is epipolar correspondence matching. The goal of this step is to propagate the edited keyframe features to all other frames by finding inter-frame correspondences. To find the correspondences, we leverage both geometric and appearance cues as follows.
@@
First, we extract visual features $\Psi_{t}$ for all frames, using intermediate outputs (i.e., the inputs to each self-attention block) across all layers in $\mathcal{E_{\operatorname{mv}}}$. Then, point correspondences across frames can be established by simply comparing their respective features at all spatial locations, and the keyframe features can be injected (based on these correspondences) into all remaining frames, avoiding expensive self-attention blocks across all frames.
@@
However, we can also leverage the fact that we can directly use 3D information to constrain the correspondence problem. In fact, we have multiple views of the scene available, and the pose and calibration matrix of cameras are known. This is without loss of generality since these assumptions are required to apply GS for scene reconstruction in the first place. Therefore, we can estimate the fundamental matrix $F$ between two views and constrain the correspondence problem to points along an epipolar line.
@@
Formally, given a feature map $\Psi_{t^{\prime}}$ corresponding to frame $I_{t^{\prime}}$, where $t^{\prime}\notin\mathcal{K}$, and the features $\Psi_{k^{*}}$ of a keyframe ${I_{k^{*}}}$, we compute a correspondence map $M_{t^{\prime}}$ as
@@
where $D$ is the cosine distance, $u$ and $v$ index the feature maps spatially, $k^{*}$ is the index of the keyframe that is the closest to frame $t^{\prime}$ (in terms of camera viewpoint), and $F$ is the fundamental matrix corresponding to the two views $t^{\prime}$ and $k^{*}$. $Fu$ is the epipolar line in view $t^{\prime}$ along which the corresponding point of $v$ in view $k^{*}$ must lie [30]. Intuitively, using the epipolar constraint significantly reduces the search space of correspondences from a plane to spatial locations along a single line. Then the correspondence is decided based on the cosine distance of the features.
In practice, we compute the correspondence maps between a frame $t^{\prime}$ and its two nearest neighbors. We then linearly combine the edited features of the corresponding keyframes to obtain $\Phi_{t^{\prime}},\forall\,t^{\prime}\in\mathcal{T}\backslash\mathcal{K}$.
@@ 4.3 Direct Reconstruction and Iterative Refinement
After obtaining a set of edited images $\mathcal{I}$ with roughly consistent content, we can directly fit $\mathcal{G}$ to the resulting images using Eq. 4, without the need for distillation. One of the notable advantages of 3D GS over NeRF is that it enables the rendering of images at full resolution and the application of robust image-level losses. Specifically, we optimize $\mathcal{G}$ using LPIPS [119] as an objective function between the edited $I^{\prime}_{t}$ and the rendered $\hat{I}_{t}$ views in Eq. 4. The LPIPS loss is insensitive to small inconsistencies in the edited images, and helps to obtain a more consistent 3D output.
Finally, we observe that even though $\mathcal{E_{\operatorname{mv}}}$ results in generally consistent edits, some inconsistencies or artifacts may remain, especially around fine and detailed textures, which in turn lead to blurry results after updating $\mathcal{G}$. We conjecture that this is due to the limited resolution of the feature space. To address this issue, after we obtain $\mathcal{G}^{\prime}$, we can again render images to be edited, repeating the process described in the previous section and re-updating the 3D model, similar to SDEdit [65] or IM-3D [64].
@@ 4.4 Partial Editing
Because 3D GS is an explicit 3D representation, if desired, a user can selectively edit a 3D scene by allowing only certain Gaussians to change, thus focusing on a specific region of interest while ensuring that the rest of the scene remains unchanged. This is a significant advantage compared to implicit radiance field representation, where there is no direct correspondence between specific parameters and regions of space, so any update to the model tends to be global.
We thus propose to first obtain a mask for the Gaussians that should undergo editing. This could be done following GaussianEditor [15] or SAGA [7]. Specifically, SAGA [7] optimize a feature field, alongside the radiance field, given user-provided seed point or mask for the object of interest.
@@ 5 Experiments
As our image editor, we use InstructPix2Pix [5], which is a fine-tuned image-to-image translation model based on Stable Diffusion [82]. We use scenes from IN2N [28] and other real-world datasets, including LLFF and Mip-NeRF360 [68, 3] to demonstrate the ability of our method on editing 3D models. We use 3D GS [43] as the 3D representation and adopt the segmentation pipeline from SAGA [7] and GaussianEditor [16] when partial editing is desired. For editing, we use 20-30 views, which are edited according to the approach outlined in Sec. 4.2 and then used to train $\mathcal{G}^{\prime}$ with 500-1000 iterations depending on the complexity of the scenes. We use LIPSP and L1 loss to train the 3D GS as suggested in IN2N [28]. We also use classifier-free guidance to control the effect strength of the editing. Most of the edits use the default setting, 7.5 for textual conditions and 1.5 for the image condition. We perform the interactive refinement every 500 iterations. We provide more detailed hyperparameters in the supplementary material.
@@
Ours GS 0.222 0.66 $\sim$ 4min
Table 1: Comparison with other editing methods. Methods based on Gaussian Splatting are much faster than NeRF-based Instruct-N2N. Besides, our DGE achieves the best performance while only using around half of the time compared to GE .
@@
We provide both qualitative and quantitative evaluations of our method. For quantitative evaluation, we followed the common practice [28, 16]. We evaluate the alignment of edited 3D models and target text prompts with CLIP similarity score, which is the cosine distance between the text and image embeddings encoded by CLIP. To be specific, we randomly sample 20 camera poses from the training dataset of the 3D models and measure the CLIP similarity scores between the rendered images and the target text prompt. Additionally, we use the CLIP directional similarity score, which is the cosine similarity distance between the image and text editing directions (target embeddings minus source embeddings), to measure the editing effect. We evaluate all the methods on a total of 3 scenes and 10 different prompts. The detailed scene and prompt pairs are provided in the supplementary.
@@ 5.1 Comparisons with Prior Work
In our experiments, we aim to demonstrate the advantages of our direct-and-consistent editing approach vs. commonly used techniques, i.e., score distillation sampling (SDS) and iterative dataset updates (IDU). We compare our method with representative text-guided 3D editing methods, IN2N [28] and GaussianEditor [15]. They both employ the IDU approach, while the former builds on a NeRF representation and the latter on Gaussian Splatting. To edit a 3D model, they both iterate between the 3D model training and image editing. Every 10 steps, they use the same 2D editor (IP2P) to edit the rendered image, replace the edited image in the 3D model training dataset, and continue the model training. Finally, we also include a baseline using the same GS framework as in our approach, but applying edits via SDS with InstructPix2Pix (IP2P + SDS). This baseline enables a direct and fair comparison between our consistent multi-view editing approach and the popular SDS-based approach.
@@
Figure 3: The comparison between with and without multi-view consistency. With the proposed multi-view consistent editing, the edited 3D GS is clear and clean while without multi-view consistent editing, it either fails to converge or leads to blurry results.
Figure 2 shows visual comparisons with the above methods. Our proposed method achieves realistic and detailed 3D editing. One notable advantage of our approach is that ensuring the consistency of the edits in image space results in 3D edits that are overall more faithful to the editing instruction (higher fidelity). In contrast, IDU and SDS work by progressively aligning the edited views of the 3D object, effectively averaging the inconsistencies from the edited images and leading to blurrier or lower-fidelity reconstructions and artifacts. For example, for the edit instruction “Make him wear a Venetian mask”, the texture of the mask generated by our method is detailed and vivid, while other methods fail to properly generate the texture. For example, GaussianEditor hardly changes the skin color, while IN2N changes the clothes and background of the 3D models as well. In the example of “Make the man look like a mosaic sculpture”, our DGE generates a consistent texture of the sculpture However, IN2N and IP2P $+$ SDS fail to converge to a reasonable editing outcome, mainly because of the inconsistency of edits during each iteration. GaussianEditor is able to generate some texture, but it is inconsistent from different views. It is worth noting that IN2N usually gets smoother results than GaussianEditor and IP2P + SDS; this is mainly due to the fact the NeRF is continuous, while GS is not.
Table 1 shows quantitative comparisons between our method and others. We achieve higher CLIP similarly score and CLIP directional similarly score, indicating that our editing 3D models align better than other methods and also follow the editing instructions.
@@ 5.2 Ablation Study
Next, we ablate the effectiveness of two main components in our editing pipeline: the multi-view consistent editing and the epipolar correspondence matching. Finally, we demonstrate the convergence speed (in terms of editing) of our approach through a qualitative example.
Figure 4: The comparison between edited 2D images with and without epipolar constraints. The one with epipolar constraints successfully matches the correspondences while the one without epipolar constraints fail to match the correspondences thus result in inconsistent multi-view editings.
Figure 5: Comparison between our DGE and GaussianEditor in terms of interaction numbers. Our method achieves realistic editing result with much fewer iterations. With iteration increases, our method also gradually refines the details.
@@
Editing without Multi-View Consistent Editing.
To validate the effectiveness of the multi-view consistent editing, we replace the multi-view consistent editing with independently editing the same number of views, and using those for fitting $\mathcal{G}^{\prime}$. Table 2 illustrates that editing with multi-view consistency is important to achieve a high CLIP score and CLIP directional score. To discuss this further, in Fig. 3, we show some intermediate 2D editing results with and without multi-view consistent editing. We observe that with multi-view consistent editing, different views get similar appearances, while per-frame editing can lead to dramatically different edits. This is especially true in the first example (“turn the man into a robot”), where the prompt allows for greater variation and occasionally even results in unsuccessful editing (third frame). As expected, fitting $\mathcal{G}^{\prime}$ to these edited views produces a low-quality 3D edited model. Our multi-view consistent approach clearly alleviates these issues. We show more detailed comparisons in supplementary materials.
Effect of the Epipolar Constraints.
In Table 2, we show that using the epipolar constraints during feature propagation (i.e. from keyframes to intermediate frames), both the CLIP score and CLIP directional score increase. Quantitatively, the difference between using or not using epipolar constraints is not large. However, we conjecture that this is mainly because the CLIP score measures the image on a high level while the epipolar constraints help to improve the detailed editing, such as textures. We better illustrate the difference in Figure 4. Without using the epipolar constraint, the resulting edits are inconsistent between the two view — red points (just below the bear’s ear). The reason is that it wrongly matches the point with another point in black which leads to failed consistent editing. The red line in the images indicates the epipolar line of the point in the red point on the keyframe. As it can only receive attention from the epipolar line, it generates the correct color.
Effect of the Number of Iterations
We show here a comparison between our method and GaussianEditor in Figure 5. The figure demonstrates that our method can rapidly edit the 3D model using only 400 iterations, while GaussianEditor requires much longer to converge. It is worth noting that the results in the column of 400 iterations are without refinement and generated within two minutes. Besides, as the interaction increases, the details of the mask rendered by the 3D model are gradually refined as we update the multi-view editing results every 500 iterations.

### v2

@@ 4.2 Consistent Multi-view Editing with Epipolar Constraint
To achieve multi-view consistent edits, we render multiple views $\mathcal{I}$ from $\mathcal{G}$ and reorder them based on their camera positions to create a relatively smooth camera trajectory, forming a video. This approach leverages recent advances in video generation and editing, and in particular, the idea of extending image generators to video without additional training [21, 94, 38, 6, 71]. Treating the sequence $\mathcal{I}$ as a video of a static scene and applying these techniques to image editors, consistency across views is analogous to temporal consistency in video editing.
Our multi-view editing process includes two sequential steps, key-view editing and feature injection. Specifically, during each denoising step, we first jointly edit select key views with spatio-temporal attention. Then, we inject the feature of the key views to other views using correspondences that are obtained from visual features guided by the epipolar constraint.
4.2.1 Key-view editing.
@@
Let $\mathcal{E}$ be a diffusion-based image editor that utilizes self-attention (e.g., SD-based). Based on the above observations, we extend $\mathcal{E}$ with a modified, spatio-temporal version of self-attention, which jointly considers all images in $\mathcal{I}$, effectively turning the single-image editor into a multi-image editor $\mathcal{E_{\operatorname{mv}}}$. Specifically, given queries $\{Q_{t}\}_{t=1}^{T}$, keys $\{K_{t}\}_{t=1}^{T}$, and values $\{V_{t}\}_{t=1}^{T}$ for each viewpoint $t$, the spatio-temporal attention block ensures that each frame attends to every other frame and is given by:
@@
$\operatorname{STAttn}(Q,K,t)=\operatorname{Softmax}\left(\frac{Q_{t}\cdot[\,K_{1},\dots,K_{T}]\ }{\sqrt{d}}\right),$ (5)
where $d$ is the embedding dimension of keys and queries. The corresponding output features for frame $t$ are then computed as $\Phi_{t}=\operatorname{STAttn}(Q,K,t)\cdot[\,V_{1},\dots,V_{T}]$ (at each layer). As the underlying image editor $\mathcal{E}$, we utilize IP2P [4], which is a common choice in the literature [25, 13, 91]. However, the attention can be computationally expensive for longer sequences. Therefore, $\mathcal{E_{\operatorname{mv}}}$ is only applied to select key views randomly selected from $\mathcal{T}=\{1,\dots,T\}$, where $T$ is the total number of images in the sequence. In practice, we sample a random subset $\mathcal{K}\subset\mathcal{T}$ at each denoising iteration. We detail the selection process in the supplement. This step results in roughly consistent edits across all key views.
@@
Figure 1: Overview. As shown on the left, our method is divided into two main parts: multi-view consistent editing with epipolar constraints and direct 3D fitting. In the multi-view editing stage, key views are randomly selected and jointly fed to the editing diffusion network to extract features with the spatial-temporal attention. To edit other frames, the features of key views are injected into the diffusion network through correspondence matching on feature maps with epipolar constraints. The detailed feature injection process is shown on the right; only features with a red border (i.e., the points following epipolar constraints) are considered for correspondence matching.
4.2.2 Feature injection.
@@
The goal of this step is to propagate the edited key view features to all other views to ensure the consistency of the edits. Specifically, we inject the features by finding inter-frame correspondences with epipolar constraints. To find the correspondences between key views and other views, we first extract visual features $\Psi_{t}$ for all views in different layers of the denoising network $\Phi$, using the intermediate outputs of $\Phi$ (i.e., the inputs to each self-attention block). Then, point correspondences across views can be established by simply comparing their respective features at all spatial locations, and the key view features can be injected (based on these correspondences) into all remaining views, avoiding expensive self-attention blocks.
@@
In addition, we can also leverage the fact that we can directly use 3D information to constrain the correspondence problem. In fact, we have multiple views of the scene available, and the pose and calibration matrix of cameras are known. This is without loss of generality since these assumptions are required to apply GS for scene reconstruction in the first place. Therefore, we can estimate the fundamental matrix $F$ between two views and constrain the correspondence problem to points along an epipolar line.
@@
Formally, given a feature map $\Psi_{t^{\prime}}$ corresponding to image $I_{t^{\prime}}$, where $t^{\prime}\notin\mathcal{K}$, and the features $\Psi_{k^{*}}$ of a keyframe ${I_{k^{*}}}$, the correspondence map $M_{t^{\prime}}$ is given by:
@@
where $D$ is the cosine distance, $u$ and $v$ index the feature maps spatially, $k^{*}$ is the index of the key view that is the closest to view $t^{\prime}$ (in terms of camera viewpoint), and $F$ is the fundamental matrix corresponding to the two views $t^{\prime}$ and $k^{*}$. $Fu$ is the epipolar line in view $k^{*}$ along which the corresponding point of $v$ in view $k^{*}$ must lie [26]. Intuitively, using the epipolar constraint significantly reduces the search space of correspondences from a plane to spatial locations along a single line. Given this constraint, the correspondence is decided based on the cosine distance of the features. In practice, we compute the correspondence maps between a view $t^{\prime}$ and its two nearest neighbors. For a given position $u$, we linearly combine the features of correspondences in the nearest two key views to obtain its edited feature.
@@ 4.3 Direct Reconstruction and Iterative Refinement
After obtaining a set of edited images $\mathcal{I}$ with roughly consistent content, we can directly fit the Gaussian mixture $\mathcal{G}$ to the resulting images using Eq. 4, without the need for distillation (SDS). One of the notable advantages of 3D GS over NeRF is that it enables the rendering of images at full resolution and the application of robust image-level losses. Specifically, we optimize $\mathcal{G}$ using LPIPS [107] as an objective function between the edited $I^{\prime}_{t}$ and the rendered $\hat{I}_{t}$ views in Eq. 4. The LPIPS loss is insensitive to small inconsistencies in the edited images, and helps to obtain a more consistent 3D output.
Finally, we observe that even though $\mathcal{E_{\operatorname{mv}}}$ results in generally consistent edits, some inconsistencies and artifacts may remain, especially around fine and detailed textures, which in turn lead to blurry results after updating $\mathcal{G}$. We conjecture that this is due to the limited resolution of the feature space. To address this issue, after obtaining $\mathcal{G}^{\prime}$, we can again render images to be edited, repeating the process described in the previous section and re-updating the 3D model, similar to SDEdit [57] or IM-3D [55].
@@ 4.4 Partial Editing
Because 3D GS is an explicit 3D representation, if desired, a user can selectively edit a 3D scene by allowing only certain Gaussians to change, thus focusing on a specific region of interest while ensuring that the rest of the scene remains unchanged. This is a significant advantage compared to implicit radiance field representations, where there is no direct correspondence between specific parameters and regions of space, so any update to the model tends to be global.
We thus propose to first obtain a mask for the Gaussians that should undergo editing. This could be done following existing works [13, 5]. In this paper, we mainly adopt the approach in [13], which segments the 2D views first and unprojects the 2D segmentation results to 3D to obtain the masks of Gaussians.
@@ 5 Experiments
As our image editor, we use InstructPix2Pix [4] for a fair comparison to prior methods using the same model. We use scenes from IN2N [25] and other real-world datasets, including LLFF and Mip-NeRF360 [60, 3] to demonstrate the ability of our method to edit 3D models. We use 3D GS [37] as the 3D representation. For editing, we use 20-30 views, which are edited according to the approach outlined in Sec. 4.2 and then train $\mathcal{G}^{\prime}$ with 500-1500 iterations depending on the complexity of the scenes. We use LPIPS and L1 loss to train the 3D GS as suggested in IN2N [25]. We also use classifier-free guidance to control the effect strength of the editing. Most of the edits use the default setting, 7.5 for textual conditions and 1.5 for the image condition. We perform the iterative refinement every 500 iterations. We provide more details in the supplementary material.
@@
Ours GS 0.226 0.67 $\sim$ 4min
Table 1: Comparison with other editing methods. Methods based on Gaussian Splatting are much faster than NeRF-based ones. Our DGE achieves the best performance at almost half the time compared to GaussianEditor.
@@
We provide both qualitative and quantitative evaluations of our method. For quantitative evaluations, we follow common practice [25, 13]. We evaluate the alignment of edited 3D models and target text prompts with CLIP similarity score, i.e., the cosine similarity between the text and image embeddings encoded by CLIP. Specifically, we randomly sample 20 camera poses from the training dataset of the 3D models and measure the CLIP similarity between the rendered images and the target text prompt. Additionally, to measure the editing effect, we use the CLIP directional similarity, i.e., the cosine similarity between the image and text editing directions (target embeddings minus source embeddings). We evaluate all methods on 3 different scenes and 10 different prompts. The detailed scene-prompt pairs are provided in the supplement.
@@ 5.1 Comparisons with Prior Work
We first compare our method with representative text-guided 3D editing methods, IN2N [25] and GaussianEditor [13]. They both employ the iterative dataset updates (IDU) approach, while the former builds on a NeRF representation and the latter on Gaussian Splatting. To edit a 3D model, they both iterate between the 3D model training and image editing. Every 10 steps, they use the same 2D editor (IP2P) to edit the rendered image, replace the edited image in the 3D model training dataset, and continue the model training. In addition, we include a baseline using the same GS framework as in our approach, but applying edits via SDS with InstructPix2Pix (IP2P + SDS). This baseline enables a direct and fair comparison between our consistent multi-view editing approach and the popular SDS-based approach. Lastly, we compare with ViCA-NeRF [16], which also focuses on editing different views consistently. Different from ViCA-NeRF, which uses 3D model depth information to project features from key views into others using a blending module, our approach incorporates the constraints from epipolar lines to avoid depth estimation errors.
In Table 1 we compare DGE to the above methods and baselines in terms of CLIP similarity score and CLIP directional similarity score. Our approach outperforms all prior work while cutting the editing time to almost half compared to GaussianEditor, the second fastest alternative.
@@
Figure 3: The comparison between with and without multi-view consistency. With the proposed multi-view consistent editing, the edited 3D GS is clear and clean, while without it, it either fails to converge or leads to blurry results.
In Figure 2 we demonstrate visual comparisons with the above methods. Our proposed method achieves realistic and detailed 3D edits. One notable advantage of our approach is that ensuring the consistency of the edits in image space results in 3D edits that are overall more faithful to the editing instruction. In contrast, IDU and SDS work by progressively aligning the edited views of the 3D object, averaging the inconsistencies from the edited images and leading to blurrier or lower-fidelity reconstructions and artifacts. For example, for the edit instruction “Make him wear a Venetian mask”, the texture of the mask generated by DGE is detailed and vivid, while other methods fail to properly generate the texture. ViCA-NeRF results in blurry edits or even fails to produce meaningful edits given prompts such as “Turn him into spider man with a mask”. It is worth noting that IN2N and ViCA-NeRF generally yield smoother results than GaussianEditor and IP2P+SDS; this is due to the fact the NeRF is continuous, while GS is not.
@@ 5.2 Ablation Study
Next, we ablate the effectiveness of two main components in our editing pipeline: the multi-view consistent editing and the epipolar correspondence matching. Finally, we demonstrate the convergence speed (in terms of editing) of our approach through a qualitative comparison with [13].
Figure 4: The comparison between edited 2D images with and without epipolar constraints. The one with epipolar constraints successfully matches the correspondences, while the other fails, thus resulting in inconsistent multi-view edits.
Figure 5: Comparison between our DGE and GaussianEditor [13] in terms of the number of iterations. Our method achieves realistic editing results with much fewer iterations. With more iterations, our method also gradually refines the details.
@@
5.2.1 Editing without Multi-View Consistency.
To demonstrate the effectiveness of the multi-view consistent editing, we replace the multi-view consistent editing, by independently editing the same views and use those to fit $\mathcal{G}^{\prime}$. Table 2 shows the importance of multi-view consistency edits, achieving higher CLIP and CLIP directional scores. To further demonstrate this, in Fig. 3 we show some intermediate 2D editing results with and without multi-view consistent editing. We observe that multi-view consistent editing produces views with similar appearances, but views can be quite different without it. This is especially true in the first example (“turn the man into a robot”), where the prompt allows for greater variation among views and may even fail to produce sensible edits for some of the frames (the third in this example). As expected, fitting $\mathcal{G}^{\prime}$ to these edited views produces a low-quality 3D edited model. Our multi-view consistent approach clearly alleviates these issues. We show more detailed comparisons in supplementary material.
5.2.2 Effect of the Epipolar Constraints.
In Table 2, we show that using the epipolar constraints during feature injection (i.e., from key views to other views), both the CLIP score and CLIP directional score increase. Quantitatively, the impact of using the epipolar constraint appears to be relatively small. However, we conjecture that this is mainly because the CLIP score assesses the image at a high level while the epipolar constraints help to improve the finer details in editing, such as textures. The importance of the epipolar constraint is more evident qualitatively. In Figure 4, the resulting edits without the epipolar constraint are inconsistent between the two views — red points (just below the bear’s ear). The reason is that it wrongly matches the point with another point in black which leads to failed consistent editing. The red line in the images indicates the epipolar line corresponding to the red point on the key view.
5.2.3 Effect of the Number of Iterations.
We additionally show a comparison between our method and GaussianEditor in Figure 5. The figure demonstrates that our method achieves high-quality edits after only 400 iterations, while GaussianEditor requires much longer for convergence. In particular, the results of the 400-iterations column do not use refinement and are generated in under two minutes. With more iterations, the details of the mask in the rendered views are gradually refined.